### v1

@@ 3.2 Discrete delay and domain decomposition
Figure 3: Left: Block linear system resulting from the spectral collocation discretisation of (3.2). The overline notation is used to indicate that the first row of the corresponding matrix is removed in order to implement boundary/continuity constraints via boundary bordering. The system is block-lower triangular, and hence one could first solve for $\bm{y}_{L}$ and then $\bm{y}_{R}$. However, this will not be the case in context of FDEs as we discuss shortly. Centre: Solution using 10- and 11-point Chebyshev grids on $[0,1/2]$ and $[1/2,1]$, respectively. Right: Convergence as $n$ is increased, where $n$ and $n+1$ are the discretisation sizes of the left and right subintervals, respectively. Using a multidomain discretisation on $[0,1/2]\cup[1/2,1]$ so that the solution is analytic on each subdomain results in geometric convergence (solid line). If only a single Chebyshev grid on $[0,1]$ is used, then algebraic convergence is observed (dashed line) due to the discontinuity in the derivative of the solution at $t=1/2$.
@@
The other two panels of Figure 3 show the computed piecewise smooth solution $\bm{y}=[\bm{y}_{L};\bm{y}_{R}]$ (centre) and the convergence of the method as the discretization size is increased (right – solid line). Since the solution is piecewise analytic on the two subintervals, geometric convergence is obtained, here using an $n$-point discretisation on the left subinterval and an ($n+1$)-point discretisation on the right. Finally, the dashed line in the third figure shows the result of applying a similar approach to discretise this DDE but without introducing the breakpoint at $t=\tfrac{1}{2}$. As expected, the jump in $y^{\prime}$ at this point results in only algebraic convergence of the numerical solution. As with most numerical methods for DDEs of this type, knowledge of the discontinuity locations is vital.
@@
The choice of using a different discretisation size on each subinterval in the example above is certainly not necessary. We do so to demonstrate that this offers no difficulty to the method. For this example, choosing an $n$-point grid on both subintervals would result in $\overline{P}(\bm{t}_{R}-\tfrac{1}{2};\bm{t}_{L})$ simplifying to an identity matrix, thus providing a sparser linear system. However, the ability to use a more general interpolation operator not only allows for differing refinement on each of the subintervals, but is necessary when subintervals are not of a uniform size, which is typically the case when there are two or more discrete delays present or the delay is nonlinear – see Example 5.
@@
$P=\left[\begin{array}[]{cccc}P_{1,1}&P_{1,2}&\cdots&P_{1,m}\\ P_{2,1}&P_{2,2}&\cdots&P_{2,m}\\ \vdots&\vdots&\ddots&\vdots\\ P_{m,1}&P_{m,2}&\cdots&P_{m,m}\\ \end{array}\right],$
@@
$[P_{j,k}]_{\ell,:}=\left\{\begin{array}[]{llll}P(\bm{[\tau}_{j}]_{\ell},\bm{t}_{k}),&T_{k}<[\bm{\tau}_{j}]_{\ell}\leq T_{k+1},\quad\text{where }\bm{\tau}_{j}=\tau(\bm{t}_{j}).\\ \bm{0}^{\top},&\rm{otherwise}.\end{array}\right.$
Note that, as in Example 3, if $\tau(t)\leq t$ then $P_{j,k}=0$ for $k>j$, i.e., $E$ is block lower-triangular. However, for FDEs this will not be the case in general.
@@
We repeat Example 3, but now solve on the domain $[0,2]$. The propagated singularities appear at $t=0.5,1,1.5$, and so we decompose $[0,2]$ into 4 uniform subintervals. The results are shown in Figure 4.
$\left[\!\!\!\begin{array}[]{ccccccccccc}\bm{e}_{1}^{\top}\\ \overline{D_{1}}+\overline{I}&\\ -\bm{e}_{n_{1}}^{\top}&\!\!\!\bm{e}_{1}^{\top}\\ \overline{P_{2,1}}&\!\!\!\overline{D_{2}}+\overline{I}\\ &-\bm{e}_{n_{2}}^{\top}&\bm{e}_{1}^{\top}\\ &\overline{P_{3,2}}&\overline{D_{3}}+\overline{I}&\\ &&-\bm{e}_{n_{3}}^{\top}&\!\!\!\bm{e}_{1}^{\top}\\ &&\overline{P_{4,3}}&\!\!\!\overline{D_{4}}+\overline{I}\end{array}\!\!\!\right]\!\left[\!\!\!\begin{array}[]{cccc}\bm{y}_{1}\\ \bm{y}_{2}\\ \bm{y}_{3}\\ \bm{y}_{4}\end{array}\!\!\!\right]=\left[\!\!\begin{array}[]{cccc}1\\ \bm{0}\\ {0}\\ \bm{0}\\ 0\\ \bm{0}\\ {0}\\ \bm{0}\end{array}\!\!\!\right]$
Refer to captionRefer to caption
Figure 4: Left: Block discretisation of the DDE on the interval $[0,2]$ with ‘breakpoints’ at $0.5,1,1.5$. Here $P_{j,k}$ is given by $P(\bm{t}_{j}-\tfrac{1}{2};\bm{t}_{k})$. Centre: Solution using 10-, 11-, 12-, 13-point grids on the respective subintervals. Right: Geometric convergence as discretisation size is increased ($n$, $n+1$, $n+2$, $n+3$-point dicretisations on each subinterval, respectively.)
Although the solution may look smooth at $t=1$ and $t=1.5$, there are discontinuities in the 2nd and 3rd derivatives of the solution at these points, respectively. However, since the solution is piecewise analytic between these points, we again observe geometric convergence in the numerical solution when using 4 subintervals with the discontinuity locations as breakpoints. Note again that the system is block lower triangular, so one can emulate the method of steps in solving first for $\bm{y}_{1}$, then $\bm{y}_{2}$, and so on.
@@ 3.4 Nonlinear delay
The approaches described above can be applied to arbitrary smooth delays functions $\tau(t)\leq t$. In such instances, locating the propagated discontinuities can be challenging, but in principle it can be done (Neves and Feldstein, 1976). With these locations known, the domain can be subdivided and multidomain approach used as described above.
@@
$y^{\prime}(t)=-y(t)-y(t^{2}-1/4),\quad y(t<0)=0,\quad y(0)=1.$ (3.5)
Here, as in the previous example, a discontinuity in the derivative of the solution is introduced at $t=1/2$, where $\tau(t)=0$. There will then be an additional discontinuity introduced in the second derivative at $t=\sqrt{3}/2$, where $\tau(t)=1/2$. In this manner, all discontinuities can be traced out, even for non-constant delays. For this example, if we solve for $0\leq t\leq 1$, these are the only two discontinuities required. (The next occurs at $t=\tfrac{1}{2}\sqrt{2\sqrt{3}+1}\approx 1.05$.) Since the subintervals $[0,1/2]$, $[1/2,\sqrt{3}/2]$, and $[\sqrt{3}/2,1]$ are of differing widths, even if the same discretisation size is used on each the resampling matrices are not trivial. We therefore discretise using the approach outlined in Section 3.2. The left panel of Figure 5 shows the result of solving with an $n=12$-point discretisation on each subinterval. The centre panel demonstrates geometric convergence as $n$ is increased. [^7] Using the method of steps, an explicit (but lengthy) solution to (3.5) can be expressed in terms of the integral of an error function. We omit the details. The final panel shows the sparsity plot of the $n=12$ discretisation.
Figure 5: Left: Analytic and computed solution to (3.5) on a 12-point grid. Centre: Geometric convergence as the discretisation size is increased. Right: Sparsity plot of the 12-point Chebyshev spectral discretisation. Here, even though the same value of $n$ is used on each subinterval, the interpolation operator is non-trivial (because the intervals are of differing widths), resulting in dense off-diagonal blocks of the discretisation.
@@ 3.5 Nonlinear problems & state-dependent delay
In spectral methods for ODEs, nonlinear problems are typically solved by a Newton–Raphson iteration, or some variant thereof–see (Driscoll and Hale, 2016, Section 4.2) for details. DDEs allow for an additional form of nonlinearity, whereby the nonlinearity occurs in the delay. This is known as state-dependent delay (Hartung et al., 2006).
Suppose a term of the form $y(g(y))$ appears in the DDE, where $g$ is some differentiable function defined on the range of $y$. To linearise, we make the following observation:
$\frac{d}{dy}\bigg{(}P(g(y))y\bigg{)}=P(g(y))+g^{\prime}(y)P^{\prime}(g(y))y=P(g(y))+g^{\prime}(y)P(g(y))\frac{dy}{dt},$ (3.6)
where we have omitted the $\bm{t}$ dependence of $P(\cdot;\bm{t})$ and the $t$ dependence of $y(t)$ for brevity. Thus the discrete Jacobian of the interpolation operator, $\frac{d}{d\bm{y}}P(g(\bm{y});\bm{t})$ , can be computed by
$\frac{d}{d\bm{y}}P(g(\bm{y}))\bm{y}=P(g(\bm{y}))+\text{diag}\Big{(}g^{\prime}(\bm{y})\circ\big{(}P(g(\bm{y}))D\bm{y}\big{)}\Big{)}.$ (3.7)
This can then be incorporated in a Newton–Raphson iteration to solve state-dependent problems.
A detailed description of the existence and uniqueness of solutions to state-dependent DDEs is beyond the scope of this paper. Furthermore, the propagation of discontinuities is significantly more complex than in the constant delay case (Neves and Feldstein, 1976). We put aside these challenges for now, and instead consider below a reasonably simple case for which a known (smooth) solution exists. A less contrived example of state-dependent delay is demonstrated in Section 5.
@@
Newton iteration:
iter  norm(res, inf)  norm(dy, inf)
@@
Refer to captionRefer to caption
Figure 6: Computed solution to the state-dependent delay problem, (3.8). Left: Quadratic convergence of the Newton iteration applied to the $n=12$ discretisation with initial guess $y=t$. Centre: Computed solution taking $n=12$. Right: Geometric convergence as $n$ is increased.
@@
A careful reader may have noticed that in the above Example we carefully chose $y(t)=\sin(t)$ so that the state-dependent delay satisfied $\tau(t,y(t))\leq t$, and was indeed delay differential equation. However, as with previous examples, this is not a requirement of the method, and a state-dependent problem with $\tau(t,y(t))\not<t$ is considered in Section 4.

### v2

@@ 3.2 Discrete delay and domain decomposition
Figure 3: Left: Block linear system resulting from the spectral collocation discretisation of (3.2). The overline notation is used to indicate that the first row of the corresponding matrix is removed in order to implement boundary/continuity constraints via boundary bordering. The system is block lower triangular, and one could first solve for $\bm{y}_{L}$ and then $\bm{y}_{R}$, akin to the method of steps. However, this will not be the case in context of FDEs as we discuss shortly. Centre: Solution using 10- and 11-point Chebyshev grids on $[0,1/2]$ and $[1/2,1]$, respectively. Right: Convergence as $n$ is increased, where $n$ and $n+1$ are the discretisation sizes of the left and right subintervals, respectively. Using a multidomain discretisation on $[0,1/2]\cup[1/2,1]$ so that the solution is analytic on each subdomain results in geometric convergence (solid line). If only a single Chebyshev grid on $[0,1]$ is used, then algebraic convergence is observed (dash-dotted line) due to the discontinuity in the derivative of the solution at $t=1/2$.
@@
The other two panels of Figure 3 show the computed piecewise smooth solution $\bm{y}=[\bm{y}_{L};\bm{y}_{R}]$ (centre) and convergence of the method as the discretisation size is increased (right, solid line), here using an $n$-point discretisation on the left subinterval and an ($n+1$)-point discretisation on the right—see Remark 1. Since the solution is piecewise analytic on the two subintervals, geometric convergence is obtained. Finally, the dashed line in the third figure shows the result of applying a similar approach to discretise this DDE but without introducing the breakpoint at $t=\tfrac{1}{2}$. As expected, the jump in $y^{\prime}$ at this point results in only algebraic convergence of the numerical solution. As with most numerical methods for DDEs of this type, knowledge of the discontinuity locations is vital (Baker et al., 1995).
@@
The choice of using a different discretisation size on each subinterval in the example above is certainly not necessary. We do so to demonstrate that this offers no difficulty to the method. For this example, choosing an $n$-point grid on both subintervals would result in $\overline{P}(\bm{t}_{R}-\tfrac{1}{2};\bm{t}_{L})$ simplifying to an identity matrix, thus providing a sparser linear system. However, the ability to use a more general interpolation operator not only allows for differing refinement on each of the subintervals, but is necessary when subintervals are not of a uniform size, which is typically the case when there are two or more discrete delays present or the delay is nonlinear—see Example 5.
@@
$P=\left[\begin{array}[]{cccc}P_{1,1}&P_{1,2}&\cdots&P_{1,m}\\ P_{2,1}&P_{2,2}&\cdots&P_{2,m}\\ \vdots&\vdots&\ddots&\vdots\\ P_{m,1}&P_{m,2}&\cdots&P_{m,m}\\ \end{array}\right],$ (3.4)
@@
$[P_{j,k}]_{\ell,:}=\left\{\begin{array}[]{llll}P(\bm{[\tau}_{j}]_{\ell},\bm{t}_{k}),&T_{k}<[\bm{\tau}_{j}]_{\ell}\leq T_{k+1},\quad\text{where }\bm{\tau}_{j}=\tau(\bm{t}_{j}).\\ \bm{0}^{\top},&\rm{otherwise}.\end{array}\right.$ (3.5)
Note that, as in Example 3, if $\tau(t)\leq t$ then $P_{j,k}=0$ for $k>j$, i.e., $P$ is block lower triangular. However, for FDEs this will not be the case in general.
@@
We repeat Example 3, but now solve on the domain $[0,2]$. The propagated singularities appear at $t=0.5,1,1.5$, and so we decompose $[0,2]$ into 4 uniform subintervals. The results are shown in Figure 4. To demonstrate that one need not use the same discretisation on each subinterval, we have arbitrarily used $n$, $n+1$, $n+2$, and $n+3$-points on each subinterval, respectively—see Remark 1.
$\left[\!\!\!\begin{array}[]{ccccccccccc}{\bm{e_{n_{1},1}}}^{\top}\\ \overline{D_{1}}+\overline{I}&\\ -{\bm{e}_{n_{1},n_{1}}}^{\top}&\!\!\!{\bm{e}_{n_{2},1}}^{\top}\\ \overline{P_{2,1}}&\!\!\!\overline{D_{2}}+\overline{I}\\ &-{\bm{e}_{n_{2},n_{2}}}^{\top}&{\bm{e}_{n_{3},1}}^{\top}\\ &\overline{P_{3,2}}&\overline{D_{3}}+\overline{I}&\\ &&-{\bm{e}_{n_{3},n_{3}}}^{\top}&\!\!\!{\bm{e}_{n_{4},1}}^{\top}\\ &&\overline{P_{4,3}}&\!\!\!\overline{D_{4}}+\overline{I}\end{array}\!\!\!\right]\!\left[\!\!\!\begin{array}[]{cccc}\bm{y}_{1}\\ \bm{y}_{2}\\ \bm{y}_{3}\\ \bm{y}_{4}\end{array}\!\!\!\right]=\left[\!\!\begin{array}[]{cccc}1\\ \bm{0}\\ {0}\\ \bm{0}\\ 0\\ \bm{0}\\ {0}\\ \bm{0}\end{array}\!\!\!\right]$
Figure 4: Left: Block discretisation of the DDE (3.2) on the interval $[0,2]$ with ‘breakpoints’ at $0.5,1,1.5$. Here $P_{j,k}$ is given by $P(\bm{t}_{j}-\tfrac{1}{2};\bm{t}_{k})$ and ${\bm{e}_{n}}_{k}$ denotes the $j$th column of the $n\times n$ identity matrix. Each nonzero entry will result in an essentially dense $n\times n$ block. Centre: Solution using 10-, 11-, 12-, 13-point grids on the respective subintervals (see Remark 1). Right: Geometric convergence as discretisation size is increased ($n$, $n+1$, $n+2$, $n+3$-point discretisations on each subinterval, respectively.)
Although the solution may look smooth at $t=1$ and $t=1.5$, there are discontinuities in the 2nd and 3rd derivatives of the solution at these points, respectively. However, since the solution is piecewise analytic between these points, we again observe geometric convergence in the numerical solution when using four subintervals with the discontinuity locations as breakpoints. Note again that the system is block lower triangular, so one can emulate the method of steps in solving first for $\bm{y}_{1}$, then $\bm{y}_{2}$, and so on.
@@ 3.3 Non-trivial history functions
Remark 3.
An important special case is periodic DDEs (PDDEs), which arise in the context of determining limit cycles of delayed dynamical systems (Michiels and Niculescu, 2007). The approach above can be modified to solve such problems by considering $\phi(t)=y({\rm mod}(\tau(t),T))$, where $T$ is the period. This is similar to the methods of (Butcher and Bobrenkov, 2011)(Borgioli et al., 2020), and (Engelborghs et al., 2000a), who use (typically low-order) polynomial-based collocation methods to solve PDDEs. In most cases, the period $T$ is then unknown, and must be solved for as part of the problem. However, the periodicity assures that the initial data will be smooth, meaning that a multidomain approach is not typically required in situations. We consider such a situation in Examples 11 and 13 below.
@@ 3.4 Nonlinear delay
The approaches described above can be applied to arbitrary smooth delays functions, $\tau(t)\leq t$. In such instances, locating the propagated discontinuities can be challenging, but in principle it can be done (Neves and Feldstein, 1976). With these locations known, the domain can be subdivided and the multidomain approach used as described above.
@@
$y^{\prime}(t)=-y(t)-y(t^{2}-\tfrac{1}{4}),\quad y(t<0)=0,\quad y(0)=1.$ (3.7)
Here, as in the previous example, a discontinuity in the derivative of the solution is introduced at $t=1/2$, where $\tau(t)=0$. There will then be an additional discontinuity introduced in the second derivative where $\tau(t)=1/2$, i.e., at $t=\sqrt{3}/2$. In this manner, all discontinuities can be traced out, even for non-constant delays. For this example, if we solve for $0\leq t\leq 1$, these are the only two discontinuities required. (The next occurs when $\tau(t)=\sqrt{3}/2$, i.e., at $t=\tfrac{1}{2}\sqrt{2\sqrt{3}+1}\approx 1.05$.) Since the subintervals $[0,1/2]$, $[1/2,\sqrt{3}/2]$, and $[\sqrt{3}/2,1]$ are of differing widths, even if the same discretisation size is used on each, the resampling matrices are not trivial. We therefore discretise using the approach outlined in Section 3.2. The left panel of Figure 5 shows the result of solving with an $n=12$-point discretisation on each subinterval. The centre panel demonstrates geometric convergence as $n$ is increased. [^7] Using the method of steps, an explicit (but lengthy) solution to (3.7) can be expressed in terms of the integral of an error function. We omit the details. The final panel shows the sparsity plot of the $n=12$ discretisation.
Figure 5: Left: Sparsity plot of the 12-point Chebyshev spectral discretisation. Here, even though the same value of $n$ is used on each subinterval, the interpolation operator is non-trivial (because the intervals are of differing widths), resulting in dense off-diagonal blocks of the discretisation. Centre: Analytic and computed solution to (3.7) on a 12-point grid. Right: Geometric convergence as the discretisation size is increased.
@@ 3.5 Nonlinear problems & state-dependent delay
In spectral methods for ODEs, nonlinear problems are typically solved by a Newton–Raphson iteration, or some variant thereof—see (Driscoll and Hale, 2016, Section 4.2) for details. DDEs allow for an additional form of nonlinearity, whereby the nonlinearity occurs in the delay. This is known as state-dependent delay (Hartung et al., 2006).
Suppose a term of the form $y(g(y))$ appears in the DDE, where $g$ is some differentiable function defined on the range of $y$. This can then be discretised using the barycentric resampling matrix on a grid $\bm{t}$ as $P(g(\bm{y});\bm{t})\bm{y}$, where $\bm{y}=y(\bm{t})$. We must linearise this expression about $\bm{y}$ to obtain the Jacobian.
To proceed, we note that, by definition,
$[\nabla\big{(}P(g(\bm{y});\bm{t})\bm{y}\big{)}]_{i,j}$ $=$ $\frac{\partial}{\partial y_{j}}[\nabla\big{(}P(g(\bm{y});\bm{t})\bm{y}\big{)}]_{i}$ (3.8)
$=$ $\frac{\partial}{\partial y_{j}}\sum_{k=1}^{n}[P(g(\bm{y});\bm{t})]_{i,k}y_{k}.$ (3.9)
Applying the product rule, we find
$[\nabla\big{(}P(g(\bm{y});\bm{t})\bm{y}\big{)}]_{i,j}$ $=$ $\sum_{k=1}^{n}\Biggl{\{}\frac{\partial}{\partial y_{j}}[P(g(\bm{y});\bm{t})]_{i,k}y_{k}+[P(g(\bm{y});\bm{t})]_{i,k}\frac{\partial}{\partial y_{j}}y_{k}\Biggr{\}}$ (3.10)
$=$ $\sum_{k=1}^{n}\Biggl{\{}\Biggl{[}\frac{\partial}{\partial y_{j}}\frac{\frac{w_{k}}{g(y_{i})-t_{k}}}{\sum_{\ell=1}^{n}\frac{w_{\ell}}{g(y_{i})-t_{\ell}}}\Biggr{]}y_{k}\Biggr{\}}+[P(g(\bm{y});\bm{t})]_{i,j}$ (3.11)
$=$ $\delta_{i,j}\Biggl{\{}\Biggl{[}\frac{d}{dy_{i}}\sum_{k=1}^{n}\frac{\frac{w_{k}}{g(y_{i})-t_{k}}}{\sum_{\ell=1}^{n}\frac{w_{\ell}}{g(y_{i})-t_{\ell}}}\Biggr{]}y_{k}\Biggr{\}}+[P(g(\bm{y});\bm{t})]_{i,j},$ (3.12)
where $\delta_{i,j}$ is the Kronecker delta. Comparing the term in braces in (3.12) to (2.1), we see that the former is precisely the derivative of the polynomial interpolant to $\{t_{k},y_{k}\}$, evaluated at $g(y_{i})$, hence
$[\nabla\big{(}P(g(\bm{y});\bm{t})\bm{y}\big{)}]_{i,j}=\delta_{i,j}\frac{d}{dy_{i}}p(g(y_{i}))+[P(g(\bm{y});\bm{t})]_{i,j}.$ (3.13)
Finally, by the chain rule, we have
$[\nabla\big{(}P(g(\bm{y});\bm{t})\bm{y}\big{)}]_{i,j}$ $=$ $\delta_{i,j}g^{\prime}(y_{i})p^{\prime}(g(y_{i}))+[P(g(\bm{y});\bm{t})]_{i,j}$ (3.14)
$=$ $\delta_{i,j}g^{\prime}(y_{i})P^{\prime}(g(y_{i});\bm{t})\bm{y}+[P(g(\bm{y});\bm{t})]_{i,j},$ (3.15)
or, in matrix form,
$\nabla\big{(}P(g(\bm{y});\bm{t})\bm{y}\big{)}=\text{diag}\Big{(}g^{\prime}(\bm{y})\cdot\big{(}P^{\prime}(g(\bm{y});\bm{t})\bm{y}\big{)}\Big{)}+P(g(\bm{y});\bm{t}),$ (3.16)
where $\cdot$ denotes pointwise multiplication. Rather than construct $P^{\prime}(g(\bm{y});\bm{t})$ explicitly, we observe that $P^{\prime}(g(\bm{y});\bm{t})=P(g(\bm{y});\bm{t})D$, and therefore that
$\nabla\big{(}P(g(\bm{y});\bm{t})\bm{y}\big{)}=\text{diag}\Big{(}g^{\prime}(\bm{y})\cdot\big{(}P(g(\bm{y});\bm{t})D\bm{y}\big{)}\Big{)}+P(g(\bm{y});\bm{t}),$ (3.17)
which can then be incorporated in a Newton–Raphson iteration to solve state-dependent problems.
A detailed description of the existence and uniqueness of solutions to state-dependent DDEs is beyond the scope of this paper. Furthermore, the propagation of discontinuities is significantly more complex than in the constant delay case (Neves and Feldstein, 1976). We put aside these challenges for now and instead consider below a reasonably simple case for which a known (smooth) solution exists. A less contrived example of state-dependent delay is demonstrated in Section 5.
@@
$k$ Residual norm Update norm
@@
Figure 6: Computed solution to the state-dependent delay problem, (3.18). Left: Quadratic convergence of the Newton iteration applied to the $n=12$ discretisation with initial guess $y=t$. Here $k$ is the iteration number and the norms are the $\ell^{\infty}$ norms on the grid points. Centre: Computed solution taking $n=12$. Right: Geometric convergence as $n$ is increased (solid) and $M_{\varepsilon}$ times the condition number of the Jacobian in the final iteration of the Newton iteration (dashed).
@@
A careful reader may have noticed that in the above example we carefully chose $y(t)=\sin(t)$ so that the state-dependent delay satisfied $\tau(t,y(t))\leq t$, and was indeed a delay differential equation. However, as with previous examples, this is not a requirement of the method, and a state-dependent problem with $\tau(t,y(t))\not<t$ is considered in Section 4.